

********************************************************************
UARS WEEKLY STATUS REPORT - 21 Dec 2001
********************************************************************
Please send future UARS Weekly Status Report submissions
to UMPGAL::STAT or "stat@umpgal.gsfc.nasa.gov"
***********
SOLSTICE
***********
SOLSTICE continues normal operations, taking solar data on all available
orbital tracking periods and stellar calibration observations whenever
possible. We continue to gather enough data to meet the SOLSTICE science
objectives during this period of reduced tape recorder operations.
Happy Winter Solstice!
***********
SUSIM
***********
SUSIM completed daily mid resolution and MgII solar scans during
periods of predicted UARS telemetry.
We are awaiting the delivery of attitude and orbit files so that
we can process SUSIM data taken after Nov. 17.
We have sent a request to the UARS FOT that the SSPP be placed in
closed loop solar pointing mode for the remainder of the UARS mission.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact
for UARS instrument operations to support re-engineering
of the UARS Mission Operations Concept to develop
capabilities for the Science Traceability Mission.
Daily processing of instrument Daily Activity Plans
continues with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 08 December 2001 (Orbit 55997, GMT Day 342) through 14
December 2001 (Orbit 56100, GMT Day 348).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
---------- ---- ----- ---------
WINDII 344/2320 56041 Power OFF
HALOE 348/1731 56098 Power OFF
The beta angle ranged between the angles of 32.72 degrees to 30.94 degrees for
this report period. The beta angle is now decreasing towards a minimum angle
of 0 degrees on 25 December 2001 (DOY 359).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 0.96 to 1.27 degrees C. Additional battery performance data for each
battery on 08 December 2001 and 14 December 2001 are provided in Appendix D.
The clock error ranged from -14.7 to 10.0 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion
of ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 10
HALOE= 1
HRDI = 1
WINDII= 1
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 4
One (1) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
* Transponder A: 00/333 - 2287.496855
* Transponder B: 01/324 - 2287.498688
GROUND SYSTEM OPERATION
Four (4) operational support problems occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations are no longer being supplied by Data Capture
Facility (DCF). This service should resume soon.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT and TAC developers continued to investigate the open DRs against the
new TACs. Several simulations were run with the simulator and the TAC as well
as internal testing of the TAC systems themselves. Although possibilities
have been excluded, the root of the problem has yet to be determined for all
but one of the three open DRs. Testing will contiue next week and these
issues will be resolved before the TACs are moved to full time operations.
The security issues have finally been resolved for getting the new Eclipse
system hooked into the Closed IONet. The systems should be configured
sometime next week and hopefully the first Noop will be sent to the simulator.
Once the Eclipse systems are connected the focus of the FOT will shift to
getting the systems tested and ready for operations.
MEETINGS
The FOT attended a UARS Eclipse Weekly on Monday, December 10, 2001.
The FOT attended a UARS Eclipse Working Meeting on Thursday, December 13,
2001.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 24 December 2001 (DOY 358). This
maneuver will be a Forward to Reverse Flight maneuver (Northern Hemisphere
viewing).
Attachments:
Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 14 December 2001
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
56000 342/0536 01-719 NBTR NBTR PBK & RTS-58
56001 342/0622 01-719 NBTR NOOPS D-342
56016 343/0715 01-720 NBTR NBTR PBK & RTS-58
56030 344/0534 01-570 NBTR NBTR PBK & RTS-58
56040 344/2141 01-721 OBC Clk Adj fr 0.00 to 20.6 msec/day
56041 344/2320 01-722 WINDII Power OFF
56045 345/0534 01-720 NBTR NBTR PBK & RTS-58
56060 346/0440 01-724 OBC/CLK Clk Adj fr 20.6 to 0.00 msec/day
56061 346/0615 01-723 NBTR PBK & RTS-58
56066 346/1452 01-723 NBTR PBK & RTS-58
56085 347/2139 01-725 OBC OBC DUMPS (NEW TAC TESTING)
56098 348/1731 01-726 HALOE Power OFF
56100 348/2057 01-727 HRDI Low Power
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 14 December 2001
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
______________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 14 December 2001
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
56040 344/2148 Clk Adj fr 0.00 to 20.6 msec/day
56060 346/0443 Clk Adj fr 20.6 to 0.00 msec/day
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 14 December 2001
08 December 2001 - GMT Day 342 Beta = 32.1 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ --------- ---------- ----- ------ -----
1 -616.0/-784.0 -3.16/-4.08 +0.0/ -0.0 10.72 0.000 0.00
2 +000.0/-134.4 +5.47/+4.48 +28.4/-14.4 25.76 1.040 20.62
3 +224.0/-224.0 +4.48/+3.82 +28.0/-14.4 25.76 1.030 20.39
14 December 2001 - GMT Day 348 Beta = 32.1 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ---------- ----- ------- -----
1 -560.0/-728.0 -3.16/-4.38 +0.0/ -0.0 10.72 0.000 0.00
2 +000.0/-504.0 +5.47/+4.48 +28.4/-12.8 26.72 1.040 18.76
3 +224.0/-112.0 +4.48/+3.82 +27.2/-12.8 26.56 1.040 18.41
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---- ---------
-no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 14 December 2001
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSY ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-075 FOT 56003 342/1040 14170 Generic Late Acquisition
01-074 NBTR 56061 346/0722 N/A NBTR at bot of Side 2, inst. rec
01-080 AP/TAC 56069 346/1913 N/A AP-7 Crashed
01-081 AP/TAC 56076 347/0713 N/A AP-1 Crashed
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-074 NBTR 56061 346/0722 N/A NBTR at bot of Side 2, inst. rec
01-080 AP/TAC 56069 346/1913 N/A AP-7 Crashed
01-081 AP/TAC 56076 347/0713 N/A AP-1 Crashed
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-75 FOT see list Dec 01 14170 Generic Late Acquisition
SEVNTFW01/50